# Is deeper always better? Replacing linear mappings with deep learning networks in the Discriminative Lexicon Model

### v1

@@ Abstract
Recently, deep learning models have increasingly been used in cognitive modelling of language. This study asks whether deep learning can help us to better understand the learning problem that needs to be solved by speakers, above and beyond linear methods. We utilise the Discriminative Lexicon Model (DLM, Baayen et al.,, 2019), which models comprehension and production with mappings between numeric form and meaning vectors. While so far, these mappings have been linear (Linear Discriminative Learning, LDL), in the present study we replace them with deep dense neural networks (Deep Discriminative Learning, DDL). We find that DDL affords more accurate mappings for large and diverse datasets from English and Dutch, but not necessarily for Estonian and Taiwan Mandarin. DDL outperforms LDL in particular for words with pseudo-morphological structure such as slend+er. Applied to average reaction times, we find that DDL is outperformed by frequency-informed linear mappings (FIL). However, DDL trained in a frequency-informed way (‘frequency-informed’ deep learning , FIDDL) substantially outperforms FIL. Finally, while linear mappings can very effectively be updated from trial-to-trial to model incremental lexical learning (Heitmeier et al., 2023b, ), deep mappings cannot do so as effectively. At present, both linear and deep mappings are informative for understanding language.
Author’s note: This is a preprint. This paper has not been peer-reviewed.
@@ 1 Introduction
Over the last decade, deep learning models have been remarkably successful, culminating in the most recent large language models (e.g. OpenAI,, 2023; Anthropic,, 2023; Gemini Team,, 2023). Increasingly, deep learning has also made its way into cognitive modelling of language with studies exploring whether deep learning architectures can model aspects of language processing. Modelling efforts have been targeted at roughly four groups of models: predicting how humans form known and novel wordforms, such as inflections (e.g. rumelhart1986learning; Cardillo et al.,, 2018; Kirov and Cotterell,, 2018; Corkery et al.,, 2019; McCurdy et al.,, 2020), using activations in deep learning models to directly predict behavioural data or brain signals (e.g. Bhatia and Richie,, 2022; Salewski et al.,, 2023; Huang et al.,, 2024; Caucheteux and King,, 2022; Schrimpf et al.,, 2021; Hale et al.,, 2022; Goldstein et al.,, 2022), using deep learning components in theoretical models of lexical processing (e.g. Seidenberg and McClelland,, 1989; Harm and Seidenberg,, 2004; Magnuson et al.,, 2020; Scharenborg,, 2008), and finally, leveraging the iterative learning ability of deep learning models to model language acquisition (e.g. Plunkett et al.,, 1992; Li et al.,, 2007).
@@
In the present paper, we address this question for the Discriminative Lexicon Model (DLM, Baayen et al.,, 2018, 2019), a model of (single) word comprehension and production. The model has been used to model the morphology of a wide range of languages (e.g. Chuang et al.,, 2020; Nieder et al.,, 2023; Heitmeier et al., 2023a, ; Chuang et al.,, 2023; van de Vijver and Uwambayinema,, 2022) but has also been successful at predicting behavioral data such as reaction times and acoustic durations (e.g. Schmitz et al.,, 2021; Stein and Plag,, 2021; Heitmeier et al., 2023b, ; Cassani et al.,, 2019; Chuang et al.,, 2021; Baayen and Smolka,, 2020; Saito,, 2023). Word forms are represented by high-dimensional binary vectors and semantics by high-dimensional real-valued vectors. In the DLM’s original formulation, comprehension is modelled with a linear mapping from word forms to word meanings and vice versa for production. Linear mappings are highly interpretable and surprisingly powerful, but also inherently limited by their linearity. In the present study we use the properties of linear and deep learning to investigated whether computational modeling can help us to better understand the learning problem that needs to be solved by speakers. Here, we are specifically interested in whether, and under what circumstances, deep learning generates insights above and beyond what more easily interpretable, linear models can tell us.
@@ 2.1 Data
In the present study we used four datasets:
@@
Estonian: This dataset was first introduced in Chuang et al., (2020). The final dataset contained 4,093 wordforms of 230 lemmas, thus on average 17.8 wordforms per lemma (range from 2 to 28). We modelled this dataset with trigrams and fasttext embeddings.
Taiwan Mandarin: This dataset was first introduced in Chuang and Baayen, (2021) based on Fon, (2004). It contained 7,349 words with one to five syllables. We used a phone-based representation where vowels are combined with a number indicating lexical tone (e.g. {CJK*}UTF8gbsn 一下子, eng. abruptly, is represented as i1.x.ia4.z.ii5) and fasttext embeddings.
@@ 2.2 Models
In the Discriminative Lexicon Model (DLM), the simplest way to represent wordforms is by first splitting wordforms into letter/phone n-grams and then marking the presence and absence of n-grams in high-dimensional binary vectors. The form vectors of all wordforms in the lexicon are stored in a form matrix $\mathbf{C}$. Word meanings are likewise represented by high-dimensional vectors, usually taken from word embedding spaces such as word2vec or fasttext (Mikolov et al.,, 2013; Grave et al.,, 2018). They are then likewise stored in a semantic matrix $\mathbf{S}$. In Linear Discriminative Learning (LDL), comprehension is then modelled via a linear mapping $\mathbf{F}$ such that
@@
$\hat{\mathbf{C}}=\mathbf{SG}.$
@@
$\hat{\mathbf{C}}$ $=g(\mathbf{S})$
Please refer to Figure 1 for an overview of the number of hidden layers and dimensions for each model in this paper. We used Rectified Linear Units (ReLUs) as activation function throughout. The comprehension models were trained with a mean squared error loss. The production models additionally had a sigmoid activation function after the final layer and were trained with a binary cross-entropy loss. We used a batchsize of 512, except for the Estonian dataset where we used a batchsize of 256. Further training details depend on the experiment and are laid out in the sections below. More details on the training procedure can be found in Heitmeier, (2024).
@@
Figure 1: DDL model architectures. The red boxes represent dense layers (matrix multiplication plus bias) with the output dimension printed inside, blue boxes represent non-linearities. Box widths reflect the input and output dimensions of the layers. Figure from Heitmeier, (2024).
@@ 3 Does DDL improve mapping accuracy?
First, we need to verify that DDL does indeed improve accuracy of the mappings compared to LDL. Here, we focus both on performance on seen data as well as unseen data. We split the data into an approximate 80/10/10 split for training/validation/test data. The data was split in such a way that all letter n-grams in the validation and test data also occurred in the training data. We trained models for all four languages for a maximum of 500 epochs, with early stopping when validation accuracy did not increase for 20 epochs (except for the BLP production model, where we waited for 50 epochs since the model did not learn at all otherwise). As a baseline, we also trained LDL models on the joint training and validation data (since LDL does not require a validation set). We evaluated all models using correlation accuracy: if the predicted vector was more correlated with its target vector than with all vectors in the remaining data, it was counted as correct. We also computed accuracy@10 for the best models for each datasets. For accuracy@10, a mapping is counted as correct, if the target vector is among the 10 most correlated vectors. The results can be found in Figure 2.
@@
It is clear that DDL outperforms LDL on both seen and unseen data for Dutch and English data where the challenge lies in mapping a diverse set of meanings. For Estonian and Mandarin on the other hand, LDL already performs surprisingly well and DDL only improves on LDL for production. However, the general increase in training accuracy across all four languages indicates that DDL does indeed afford higher mapping precision.
@@
Next, we investigated for which words DDL outperforms LDL. We hypothesised that DDL would perform better than LDL for words which are harder to distinguish from other words. These are words with n-grams that are shared with many other words, shorter words and words in a dense semantic neighbourhood. We used these predictors (plus word frequency as a control) to predict the difference in correlation between predicted and target vectors in LDL and DDL using a Generalised Additive Model (GAM, Wood,, 2011; more details on how we operationalised these concepts can be found in Heitmeier,, 2024). Figure 3 shows that DDL indeed outperforms LDL for words with higher cue overlap (i.e. a word’s cues occur in many other words), shorter words and words in a denser orthographic neighbourhood (higher Coltheart’s N, Coltheart,, 1977). Moreover, it performs better on words with higher frequency. While semantic density (measured as the correlation between the target semantic vector and its closest semantic neighbour) seems to have a U-shaped effect, this is due to collinearity with the other predictors (Friedman and Wall,, 2005): in a GAM with only semantic density as a predictor, the measure shows the expected effect of an advantage for DDL for words in denser semantic neighbourhood.
@@
Figure 3: Measures of between-word similarity predicting the difference between target correlation in DDL and LDL (target correlation DDL - target correlation LDL). For higher values of correlation difference, DDL outperforms LDL. We find that in general, DDL outperforms LDL for words that are similar to other words (i.e. words with higher cue overlap, a dense orthographic neighbourhood, shorter words and more frequent words). Figure from Heitmeier, (2024).
@@
We also qualitatively examined words that showed particularly stark differences in correlation, favoring either LDL or DDL. The results can be found in Table 2. LDL outperforms DDL if words contain unique cues such as px or tv. DDL on the other hand shows a clear advantage for words which contain suffixes expressing a range of different meanings such as slender, barriers or choler which all contain trigrams which are often used in comparatives. Consequently, LDL places them in the semantic neighbourhood of other comparatives, while DDL maps them to their correct neighbourhood. This ability of DDL to work around the re-use of trigrams to express different meanings points to an interesting issue: language is full of re-use of sublexical features such as the -s ending expressing plurality (see also Shafaei-Bajestan et al.,, 2022; Plag et al.,, 2015), 3rd person singular, possessive or simply being a part of a word’s stem, or the -er ending which can express as varied meanings as comparative (e.g. hungrier), agent (e.g. teacher), instrument (e.g. poker), something associated with an action (e.g. diner) or denoting the origin of a person (e.g. Londoner) (Plag,, 2003, p.89). Any system modelling morphology needs to be able to deal with this “bricolage” (Lévi-Strauss,, 1962). Apparently, DDL is much more effective at doing so than LDL. This does not seem to be just an issue of precision: both Table 2 and the accuracies@10 in Figure 2 suggest that for LDL, the correct target word does not even make it to the top 10 most correlated words. A replication of this analysis for Dutch can be found in the Supplementary Materials.
@@
by by, standby, bye, written, thereby, brigades, dismissed, copied, outplayed, routed chastised, cajoled, bypassed, imposed, enforced, censured, challenged, embraced, shamed, entrenched -0.47
my my, her, me, your, own, myself, his, friend, our, dearest my, fathers, bruised, dearest, dad, childhood, dear, treasured, scarred, friend -0.41
px px, width, height, border, size, square, margin, padding, serif, background px, font, width, serif, size, apex, height, margin, cornered, tone -0.41
foursquare foursquare, twitter, mobile, hangouts, badges, twitted, yelp, badge, tweet, pollards foursquare, taproom, tavern, hangouts, beard, goatee, lapel, moustache, alehouse, stubble -0.35
tv tv, telly, cable, broadcasts, reruns, channels, airing, tuner, channel, viewers tv, telly, martin, shaw, john, cable, tom, mike, lightbulb, michael -0.34
@@
slang slashing, chopping, mangling, slamming, bashing, slating, slanting, hewing, chiding, slantwise slang, lingo, slangs, idiom, slur, patois, spelling, idioms, accent, rhymes 0.58
slender slighter, sturdier, trendier, rougher, livelier, finer, stouter, chunkier, bulkier, heartier slender, lithe, graceful, sinuous, lanky, tendrils, curving, threadlike, arched, leafless 0.58
charming arming, baiting, goading, harming, snaring, singeing, luring, robbing, maiming, gassing charming, quaint, demure, charm, classy, genteel, homely, urbane, serene, chatty 0.59
barriers livelier, chattier, sturdier, flashier, rowdier, trendier, sketchier, ruder, heartier, fleshier barriers, barrier, roadblocks, hurdle, curbs, impede, hindrance, crossings, hurdles, gaps 0.59
choler chunkier, wiser, moister, heartier, blunter, fattier, gunsmith, braver, plainer, trimmer choler, wherefore, foulness, rancour, straightway, withal, chequer, thine, vileness, smote 0.63
@@
Table 2: Words for which LDL outperforms DDL the most (a) and for which DDL outperforms LDL the most. The two “semantic neighbours” columns show the words of the 10 semantic vectors which are most correlated with the vectors predicted by DDL and LDL respectively. The rightmost column shows the difference in target correlation between LDL and DDL. Table reproduced from Heitmeier, (2024).

### v2

@@ Abstract
Recently, deep learning models have increasingly been used in cognitive modelling of language. This study asks whether deep learning can help us to better understand the learning problem that needs to be solved by speakers, above and beyond linear methods. We utilise the Discriminative Lexicon Model introduced by Baayen and colleagues, which models comprehension and production with mappings between numeric form and meaning vectors. While so far, these mappings have been linear (Linear Discriminative Learning, LDL), in the present study we replace them with deep dense neural networks (Deep Discriminative Learning, DDL). We find that DDL affords more accurate mappings for large and diverse datasets from English and Dutch, but not necessarily for Estonian and Taiwan Mandarin. DDL outperforms LDL in particular for words with pseudo-morphological structure such as chol+er. Applied to average reaction times, we find that DDL is outperformed by frequency-informed linear mappings (FIL). However, DDL trained in a frequency-informed way (‘frequency-informed’ deep learning, FIDDL) substantially outperforms FIL. Finally, while linear mappings can very effectively be updated from trial-to-trial to model incremental lexical learning, deep mappings cannot do so as effectively. At present, both linear and deep mappings are informative for understanding language.
@@ 1 Introduction
Over the last decade, deep learning models [^1] Throughout this paper, we use the term ‘deep learning’ to refer to neural networks with at least one hidden layer (with a non-linear activation function). While it can be argued whether a neural network with a single hidden layer can indeed be called ‘deep’, the definitions of deep learning models usually only require such models to have “multiple (non-linear) processing layers” (e.g. Arnold et al.,, 2011; LeCun et al.,, 2015) and for reasons of simplicity we therefore include models with a single non-linear hidden layer in the definition. Crucially, and contrary to models without any hidden layers and no activation functions, such models are able to learn non-linear functions (e.g. Webb and Lowe,, 1990; Widrow and Lehr,, 1990). have been remarkably successful, culminating in the most recent large language models (e.g. OpenAI,, 2023; Anthropic,, 2023; Gemini Team,, 2023). Increasingly, deep learning has also made its way into cognitive modelling of language with studies exploring whether deep learning architectures can model aspects of language processing. Modelling efforts have been targeted at roughly four groups of models: predicting how humans form known and novel wordforms, such as inflections (e.g. Rumelhart and McClelland,, 1986; Cardillo et al.,, 2018; Kirov and Cotterell,, 2018; Corkery et al.,, 2019; McCurdy et al.,, 2020), using activations in deep learning models to directly predict behavioural data or brain signals (e.g. Bhatia and Richie,, 2022; Salewski et al.,, 2023; Huang et al.,, 2024; Caucheteux and King,, 2022; Schrimpf et al.,, 2021; Hale et al.,, 2022; Goldstein et al.,, 2022), using deep learning components in theoretical models of lexical processing (e.g. Seidenberg and McClelland,, 1989; Harm and Seidenberg,, 2004; Magnuson et al.,, 2020; Scharenborg,, 2008), and finally, leveraging the iterative learning ability of deep learning models to model language acquisition (e.g. Plunkett et al.,, 1992; Li et al.,, 2007).
@@
In the present paper, we address this question for the Discriminative Lexicon Model (DLM, Baayen et al.,, 2018, 2019), a model of (single) word comprehension and production. The model has been used to model the morphology of a wide range of languages (e.g. Chuang et al.,, 2020; Nieder et al.,, 2023; Heitmeier et al., 2023a, ; Chuang et al.,, 2023; van de Vijverand Uwambayinema,, 2022) but has also been successful at predicting behavioral data such as reaction times and acoustic durations (e.g. Gahl and Baayen,, 2024; Stein and Plag,, 2021; Heitmeier et al., 2023b, ; Cassani et al.,, 2019; Chuang et al.,, 2021; Baayen and Smolka,, 2020; Saito,, 2023). Word forms are represented by high-dimensional binary vectors and semantics by high-dimensional real-valued vectors. In the DLM’s original formulation, comprehension is modelled with a linear mapping from word forms to word meanings and vice versa for production. Linear mappings are highly interpretable and surprisingly powerful, but also inherently limited by their linearity. In the present study we use the properties of linear and deep learning to investigate whether computational modeling can help us to better understand the learning problem that needs to be solved by speakers. Since we investigate single-word learning rather than word sequences and we believe that in the context of cognitive modelling the analysability of models is very important, we focus on using simple dense neural networks rather than more complex deep learning architectures such as Recurrent Neural Networks (RNNs) or transformers. We are specifically interested in whether, and under what circumstances, deep learning generates insights above and beyond what more easily interpretable, linear models can tell us.
@@ 2.1 Data
In the present study we used four datasets [^2] Data, code and analysis scripts can be found in the Supplementary Materials at https://osf.io/tyfdr/).:
@@
Estonian: This dataset was first introduced in Chuang et al., (2020). The final dataset contained 4,093 wordforms of 230 lemmas, thus on average 17.8 wordforms per lemma (range from 2 to 28). We modelled this dataset with trigrams and fasttext embeddings (Grave et al.,, 2018).
Taiwan Mandarin: This dataset was first introduced in Chuang and Baayen, (2021) based on Fon, (2004). It contained 7,349 words with one to five syllables. We used a phone-based representation where vowels are combined with a number indicating lexical tone (e.g. 一下子, eng. ‘abruptly’, is represented as i1.x.ia4.z.ii5) and fasttext embeddings (Grave et al.,, 2018).
@@ 2.2 Models
In the Discriminative Lexicon Model (DLM) comprehension is modelled as a mapping from a word’s form to its meaning, and production as a mapping from meaning to form. The simplest way to represent wordforms is by first splitting wordforms into letter/phone n-grams and then marking the presence and absence of n-grams in high-dimensional binary vectors. The form vectors of all wordforms in the lexicon are stored in a form matrix $\mathbf{C}$. Word meanings are likewise represented by high-dimensional vectors, usually taken from pre-trained word embedding spaces such as word2vec or fasttext (Mikolov et al.,, 2013; Grave et al.,, 2018). Details on which embedding vectors were used for which dataset can be found in Section 2.1. The embedding vectors are then likewise stored in a semantic matrix $\mathbf{S}$. In Linear Discriminative Learning (LDL), comprehension is then modelled via a linear mapping $\mathbf{F}$ such that
@@
$\hat{\mathbf{C}}=\mathbf{SG}$
with $\hat{\mathbf{S}}$ and $\hat{\mathbf{C}}$ indicating predicted $\mathbf{S}$ and $\mathbf{C}$ matrices respectively, borrowing notation from statistics. There are several methods available for calculating $\mathbf{F}$ and $\mathbf{G}$, such as standard least squares regression which optimises across all word forms equally, or “frequency-informed learning” which uses a form of weighted regression to inform the mappings about word forms’ frequencies (Heitmeier et al., 2023a, ). We used various methods throughout this study and give more detail in the relevant sections.
Once $\mathbf{F}$ and $\mathbf{G}$ are in place, they can be used to predict semantic and form vectors for seen or unseen word forms $w$ via
$\hat{\mathbf{s}}_{w}=\mathbf{c}_{w}\mathbf{F}$
$\hat{\mathbf{c}}_{w}=\mathbf{s}_{w}\mathbf{G}$
where $\mathbf{c}_{w}$ and $\mathbf{s}_{w}$ are the form and meaning vectors of $w$ and $\hat{\mathbf{s}}_{w}$ and $\hat{\mathbf{c}}_{w}$ the model’s predicted vectors. A thorough introduction to LDL can be found in Heitmeier et al., 2025a .
@@
$\hat{\mathbf{C}}$ $=g(\mathbf{S}).$
Since multi-layer neural networks can in principle approximate any mathematical function (with some limitations, Hornik et al.,, 1989), we denote them with $f$ and $g$.
Please refer to Figure 1 for an overview of the number of hidden layers and dimensions for each model in this paper (one model per dataset and comprehension/production). We used Rectified Linear Units (ReLUs) [^4] Due to space limitations, we refrain here from defining terms commonly used in machine learning literature. A more accessible introduction to Deep Discriminative Learning including term definitions can be found in Chapter 11 of Heitmeier et al., 2025a . as activation function throughout. The comprehension models were trained with a mean squared error loss. For production models, we opted to regard the prediction of which n-grams are part of a wordform as a multi-class multi-label classification problem where all n-grams are “classes” and multiple n-grams can be selected as “labels”. As commonly done in multi-class multi-label classification models, we therefore added a sigmoid function after the final layer of the production models to restrict the final predicted values between 0 and 1 (approximating probabilities) and trained them with a binary cross-entropy loss. Note that adding a sigmoid + binary cross-entropy loss to the model in this way does not add any additional trainable parameters. We used a batchsize of 512, except for the Estonian dataset where we used a batchsize of 256. Unless indicated otherwise, we used the Adam optimizer (Kingma,, 2014) with a learning rate of 0.001. Further justification for these modelling choices can be found in Heitmeier et al., 2025b . Other training details depend on the experiment and are laid out in the sections below.
We note here that we do not regard the different architectures across the different datasets and tasks as being informative about the cognition of speakers of the respective languages. Instead, we think that the architectures are determined by the datasets. The datasets vary considerably in size which influences the networks’ required sizes. The languages’ structures presumably also influence the architectures, but even regarding this point, we are wary of making any claims about speakers’ cognition given that the kind of neural networks we are using are not particularly biologically plausible and how words were selected varies considerably across datasets. Instead, we view the presented architectures as the optimal models for solving the specific mapping problems for the specific datasets (within the scope of hyperparameters tested). As such, we are interested in what can theoretically be learned by a statistical system and testing whether what is learned by such a system allows insights about behaviour as well.
@@
Figure 1: DDL model architectures. The red boxes represent dense layers (matrix multiplication plus bias) with the output dimension printed inside, blue boxes represent non-linearities. Box widths reflect the input and output dimensions of the layers. Figure from Heitmeier et al., 2025b .
@@ 3 Does DDL improve mapping accuracy?
First, we need to verify that DDL does indeed improve accuracy of the mappings compared to LDL. Here, we focus both on performance on seen data as well as unseen data. We split the data into an approximate 80/10/10 split for training/validation/test data. The data was split in such a way that all letter n-grams in the validation and test data also occurred in the training data. We trained models for all four languages for a maximum of 500 epochs, with early stopping when validation accuracy did not increase for 20 epochs for the comprehension and 50 epochs for the production models. As a baseline, we also trained LDL models on the joint training and validation data (since LDL does not require a validation set) [^5] Even though DDL is not directly trained on the validation data, it is nevertheless used in training to select the epoch after which to stop training when performance on the validation data is best. Thus, even though the validation data is not directly used to calculate the loss and update the model’s weights, it nevertheless influences the trained model. Due to this crucial role in the training of the DDL models, it is important to train the linear models on both the training and validation data.. We evaluated all models using correlation accuracy: if the predicted vector was more correlated with its target vector than with all vectors in the remaining data, it was counted as correct. We also computed accuracy@10 for the best models for each datasets. For accuracy@10, a mapping is counted as correct, if the target vector is among the 10 most correlated vectors. The results can be found in Figure 2.
@@
It is clear that DDL outperforms LDL for comprehension only for the English data where the challenge lies in mapping a diverse set of meanings. For Estonian and Mandarin on the other hand, LDL already performs surprisingly well and DDL only improves on LDL for production. However, DDL outperforms LDL for all datasets in production and the general increase in training accuracy across all four languages indicates that DDL does indeed afford higher mapping precision.
@@
Next, we investigated for which words DDL outperforms LDL. We hypothesised that DDL would perform better than LDL for words which are harder to distinguish from other words. These are words with n-grams that are shared with many other words, shorter words and words in a dense semantic neighbourhood. We used these predictors (plus word frequency as a control) to predict the difference in correlation between predicted and target vectors in LDL and DDL using a Generalised Additive Model (GAM, Wood,, 2011; more details on how we operationalised these concepts can be found in Heitmeier et al., 2025b, ). Figure 3 shows that DDL indeed outperforms LDL for words with higher cue overlap (i.e. a word’s cues occur in many other words), shorter words and words in a denser orthographic neighbourhood (higher Coltheart’s N, Coltheart,, 1977). Moreover, it performs better on words with higher frequency. While semantic density (measured as the correlation between the target semantic vector and its closest semantic neighbour) seems to have a U-shaped effect, this is due to collinearity with the other predictors (Friedman and Wall,, 2005): in a GAM with only semantic density as a predictor, the measure shows the expected effect of an advantage for DDL for words in denser semantic neighbourhood.
@@
Figure 3: Measures of between-word similarity predicting the difference between target correlation in DDL and LDL (target correlation DDL - target correlation LDL). For higher values of correlation difference, DDL outperforms LDL. The lower x-axes show the centered and scaled values as entered into the GAMs; the upper x-axes show the original values for reference. We find that in general, DDL outperforms LDL for words that are similar to other words (i.e. words with higher cue overlap, a dense orthographic neighbourhood, shorter words and more frequent words). Figure adapted from Heitmeier et al., 2025b .
@@
We also qualitatively examined words that showed particularly stark differences in correlation, favoring either LDL or DDL. The results can be found in Table 2. LDL outperforms DDL if words contain unique cues such as px or oz. DDL on the other hand shows a clear advantage for words which contain suffixes expressing a range of different meanings such as barriers or choler which all contain trigrams which are often used in comparatives. Consequently, LDL places them in the semantic neighbourhood of other comparatives, while DDL maps them to their correct neighbourhood. This ability of DDL to work around the re-use of trigrams to express different meanings points to an interesting issue: language is full of re-use of sub-lexical features, such as the -s ending, which can express plurality (see also Shafaei-Bajestan et al.,, 2022; Plag et al.,, 2015), third person singular, or possessive or simply be part of a word’s stem, or the -er ending, which can express varied meanings such as comparative (e.g., hungrier), agent (e.g., teacher), instrument (e.g., poker), or something associated with an action (e.g., diner) or can denote the origin of a person (e.g., Londoner) (Plag,, 2003, p.89). Any system modelling morphology needs to be able to deal with this “bricolage” (Lévi-Strauss,, 1962). Apparently, DDL is much more effective at doing so than LDL. This does not seem to be just an issue of precision: both Table 2 and the accuracies@10 in Figure 2 suggest that for LDL, the correct target word does not even make it to the top 10 most correlated words. A replication of this analysis for Dutch can be found in the Supplementary Materials.
@@
by by, standby, bye, written, thereby, brigades, dismissed, copied, outplayed, routed bypassed, policed, weeded, cleared, entrenched, enforced, watered, encroached, deterred, imposed -0.51
px px, width, height, border, size, square, margin, padding, serif, background px, margin, apex, equals, border, serif, font, zero, cornered, width -0.50
my my, her, me, your, own, myself, his, friend, our, dearest my, friend, dearest, dear, proudest, thankful, hearts, beloved, fathers, coolest -0.44
if if, unless, sure, else, whether, because, assume, not, does, should if, unless, assume, sure, whereas, therefore, neither, likely, maybe, suppose -0.40
oz oz, ounce, ounces, bottle, quart, gallon, jar, servings, cologne, spray oz, ounce, ounces, bottle, juice, gallon, shampoo, tumbler, cleanser, pint -0.38
@@
lungs lunches, lunge, slogs, dragged, luncheons, lunges, lurches, clamoured, jaunts, dives lungs, lung, chest, breath, windpipe, coughing, exhale, larynx, breathe, throat 0.54
scarves carves, carving, starves, carvers, carver, corpses, gouges, carved, sews, sculpts scarves, scarfs, sweaters, stoles, blankets, hats, headbands, afghans, robes, capes 0.55
leans leaches, cleanses, scoundrels, deign, nazis, cleans, traitors, betters, purveys, kikes leans, pulls, mutters, nods, pushes, straightens, crouches, grabs, squeezes, smirks 0.56
barriers livelier, chattier, sturdier, flashier, rowdier, trendier, sketchier, ruder, heartier, fleshier barriers, barrier, impede, roadblocks, blocking, curbs, hurdle, shackles, hinder, doorways 0.56
choler chunkier, wiser, moister, heartier, blunter, fattier, gunsmith, braver, plainer, trimmer choler, slackness, rancour, wherefore, quinsy, foulness, rapine, glumness, stoutness, largeness 0.59
@@
Table 2: Words for which LDL outperforms DDL the most (a) and for which DDL outperforms LDL the most (b). The two “semantic neighbours” columns show the words of the 10 semantic vectors which are most correlated with the vectors predicted by LDL and DDL respectively. The rightmost column shows the difference in target correlation between LDL and DDL. LDL outperforms DDL for words with unique n-grams, while DDL outperforms LDL for words with endings which can express various meanings (e.g. the -er ending). In the latter case, LDL tends to place words in the semantic neighbour of words with the same ending, while DDL puts them in the neighbourhood of words with similar meanings.